QD,HC= 8,000 5PHC

QS,HC= 10PCH 2,500

  1. Show on a graph the current equilibrium price (PHC*) and the current equilibrium quantity (QHC*). Label this point “A”. (PHC on the vertical axis and QHC on the horizontal axis). Do not forget to include the value of the intercepts on the vertical and horizontal axes. Make sure to properly label the curves and the axes. MAKE SURE TO SHOW ALL YOUR WORK.
  2. Now suppose that the government introduces a subsidize given to health care providers of $200 per service rendered. Given this scenario, calculate the new price that patients are paying out of pocket per service (Pb), the new price health care providers are receiving per service (PS), and the new equilibrium quantity (QHC,NEW*). Make sure to show all your work.
Answer #1

CHOC go: 05 eam: HC HC =8000 - 5 PHC = 1o Prie - 2500 (a at 1600k s 2 8000 - 5Pue = 10PHC-2500 10500 = 15 Pic : 4 = 0 = 7000